OpenStudy (anonymous):

express the sum of terms of 4√-12 and 3√-27 in simplest radical form in terms of i

OpenStudy (anonymous):

\[8i \sqrt{3} + 9i \sqrt{3} = 17i \sqrt{3}\]\[4\sqrt{-12} = 8i \sqrt{3}\]\[3\sqrt{-27} = 9i \sqrt{3}\]

OpenStudy (anonymous):

You put what is given into simplist form and then add them together. Thats what the first line shows.

OpenStudy (anonymous):

thanks